# Let’s Measure Information Step-by-Step: LLM-Based Evaluation Beyond Vibes

### v1

@@ 3.4 Fundamental Limits
Let $B$ be any distribution-free estimator providing a $(1-\delta)$ confidence lower bound on $I_{f}(X;Y)$, derived from a finite sample empirical type $\mathcal{T}(S)$ of size $N$. For integers $k\geq 1$ and $N\geq 2$, with probability at least $1-\delta-1/k$ over samples $S\sim P_{XY}^{N}$:
$B\bigl{(}\mathcal{T}(S),\delta\bigr{)}\leq\frac{1}{2kN^{2}}f(2kN^{2})+\left(1-\frac{1}{2kN^{2}}\right)f(0).$
@@
Figure 1 (Right) shows the adversarial “mode collapse” construction that drives the bound: keep the largest $kN^{2}$ parts of the response distribution unchanged, spread the next $kN^{2}$ likely responses uniformly at height $1/(2kN^{2})$, and drop the rest. Because this surrogate distribution has only $2kN^{2}$ support points, Lemma D.1 forces its $f$-information below $\frac{1}{2kN^{2}}f(2kN^{2})+\bigl{(}1-\frac{1}{2kN^{2}}\bigr{)}f(0)$. This is the dashed level in the figure.
Each orange atom has mass at most $1/(kN^{2})$, so a union bound over the $\binom{N}{2}$ possible duplicate pairs shows a pure sample (no tail element repeats) occurs with probability $\geq 1-\tfrac{1}{2k}$. On every pure sample the estimator’s $(1-\delta)$ guarantee keeps $B$ below the dashed ceiling, hence $\Pr\bigl{[}B(\mathcal{T}(S),\delta)>\text{ceiling}\bigr{]}\leq\delta+\tfrac{1}{k}$, which rearranges to the claimed bound. ∎
This theorem reveals how different choices of the convex function $f$ affect achievable estimation accuracy. For unbounded measures such as Shannon mutual information where $f(t)=\log t$, the bound grows logarithmically with sample size, requiring exponentially large samples for meaningful guarantees. Bounded measures exhibit polynomial decay, enabling practical implementation with finite samples.
These limits extend the estimation bounds first studied by McAllester and Stratos (2020) to the general class of $f$-divergences. The theorem makes a clear empirical prediction: bounded measures (like TVD where $f$ is bounded) should maintain robustness under adversarial manipulation, while unbounded measures (like Shannon MI where $f(t)=t\log t$) should degrade catastrophically. Section 5 validates this prediction directly. TVD-MI degrades by only 31.2% while MI/DoE degrades by 136.7% under identical attacks, confirming the practical importance of choosing bounded information measures.
@@ 4 Study Setup
We designed and pre-registered  [^1] Pre-registration available at https://osf.io/c7pum  [^2] Code available at https://github.com/zrobertson466920/llm-peer-prediction/tree/main an evaluation study to test whether information-theoretic mechanisms can reliably detect strategic manipulation in AI-generated content across diverse domains. Our study addresses three primary research questions:
RQ1: Can information-theoretic mechanisms detect strategic manipulation? We hypothesized that mechanisms grounded in information theory would distinguish information-preserving agents from those that degrade or manipulate content (H1a: effect size d > 0.5).
RQ2: How does task compression affect mechanism performance? We predicted linear degradation with increasing compression ratios (H1b), expecting the most compressed tasks to show the weakest discrimination.
RQ3: What properties make mechanisms robust? We hypothesized that bounded mechanisms like TVD-MI would show superior stability (H1c) and resistance to gaming (H2c).
Our pre-registered hypotheses yielded unexpected results. While all information-theoretic mechanisms achieved universal success (supporting H1a), they exhibited an inverted-U relationship with compression rather than linear degradation. This prompted investigation into the information structure underlying mechanism performance.
4.1 Minimizing Confounders in the Design
Domain Selection. We selected 10 domains spanning compression ratios (output length : input length) from 1.1:1 (translation) to 20.2:1 (peer review), covering task types (translation, summarization, peer review) and difficulty levels. This range allows us to isolate effects of information compression from task-specific artifacts.
Agent Strategy Taxonomy. We developed a taxonomy of agent behaviors grouped into two primary categories based on their approach to the task:
Good Faith Agents (information-preserving) attempt to complete the task helpfully, though potentially with stylistic biases:
  • Faithful: Direct, accurate task completion (4-6 variants per domain)
  • Style: Alternative presentation while preserving information (10-16 variants per domain)
Problematic Agents (information-degrading) either manipulate content or provide low-quality outputs:
  • Strategic: Deliberate information distortion or misrepresentation (4-10 variants per domain)
  • Low Effort: Minimal effort or generic responses (4-5 variants per domain)
Each domain included 29-30 agent conditions, carefully balanced across categories. We used an LLM (Claude-4 Opus) to generate the specific prompts for each task type. Summarization tasks had keywords swapped out to maintain consistency. For example, in summarization, “Fact Manipulation” (Strategic) alters key details while “Ultra Concise” (Low Effort) provides minimal information. This taxonomy, while containing some subjective boundary cases, provides a principled framework for testing discrimination capabilities. The complete taxonomy includes 30 conditions for translation, 29 for summarization, and 30 for peer review (see Appendix for full details).
Controlled Implementation. All agent responses were generated using GPT-4o-mini with temperature 0.7 for consistency across conditions. We employed identical prompting structures across domains, varying only the task-specific instructions. This standardization isolates the effects of agent strategies from implementation artifacts.
Comprehensive Evaluation. We computed all pairwise comparisons between agent conditions bidirectionally, yielding 870 comparisons per example. This approach provides statistical power while revealing patterns in mechanism behavior.
4.2 Key Methods
Mechanism Implementations. We implemented four primary evaluation mechanisms representing different theoretical approaches.
Computational Cost Comparison. All mechanisms require equivalent computational resources for practical deployment. In practice each mechanism requires only a single API call per evaluation similar to LLM judges. The key difference lies in what question we ask ("do these share information?" vs "which is better?"), not computational complexity.
Mutual Information (MI/DoE). We estimate mutual information using the Difference of Entropies approach:
$\text{MI}(A;B)=H(A)-H(A|B)$ (7)
We compute this using log probabilities from Llama 3.3-70B-Instruct, chosen for its accessibility and reasonable performance. The normalized log probability serves as our entropy estimate. This approach is particularly efficient as it requires only token log-probabilities without full generation, making it faster and cheaper than streaming-based methods.
Generative Peer Prediction Mechanism (GPPM). We implement GPPM as our primary baseline following Lu et al. (2024):
$\text{GPPM}(A,B)=\log P(A|B)+\log P(B|A)$ (8)
This mechanism lacks the unique characterization of $f$-MI but provides a principled comparison point from the peer prediction literature. Like MI/DoE, GPPM requires only log-probability computation without token generation and offers computational efficiency comparable to or better than standard evaluation approaches.
Total Variation Distance MI (TVD-MI). Our implementation uses an LLM critic (GPT-4o-mini) to estimate bounded mutual information. The critic evaluates whether two responses show evidence of addressing the same source material and outputs categorical judgments of whether the responses share a source. This categorical approach provides the lower bound property for gaming resistance. See Figure 1 (Left Panel 3) for an illustration.
LLM Judge. As a standard baseline, we implement pairwise comparison using GPT-4o-mini in two modes:
  • With Context: Judge sees the original input (article/paper)
  • Without Context: Judge evaluates responses in isolation
This tests whether access to ground truth affects discrimination ability.
Statistical Analyses. For each domain, we computed Cohen’s d effect sizes comparing Good Faith versus Problematic agent categories, with 95% confidence intervals via bootstrap resampling (1000 iterations). We employ bootstrap resampling over individual pairwise comparisons to compute effect sizes (Cohen’s $d$) between Good Faith and Problematic categories. We compute 95% confidence intervals using 1000 bootstrap iterations. We then perform correlation analysis between mechanisms and reference-based baseline metrics (BLEU/ROUGE/LLM Judge) Papineni et al. (2002); Lin (2004); Zheng et al. (2023).
To investigate compression, we initially fit linear models as pre-registered. Following unexpected results, we extended our analysis to include quadratic models and measured fit improvement. For each domain, we computed the effective rank of the TVD-MI score matrix across agent conditions as $\frac{\bigl{(}\sum_{i}|\lambda_{i}|\bigr{)}^{2}}{\sum_{i}\lambda_{i}^{2}}$ where $\lambda_{i}$} are the eigenvalues of the matrix. This spectral definition quantifies the effective dimensionality of the agents’ response space and measures information diversity within each domain.
@@ 4.2.1 Measurement Tampering Experiments
Our measurement tampering experiments directly test the adversarial construction from Theorem 3.5. The theoretical proof shows how adversaries can degrade estimation by redistributing probability mass from tail elements to create more uniform distributions. We implement practical analogues of this theoretical attack:
@@
    Format standardization: Replaces diverse formatting with consistent structure, analogous to the theoretical construction’s redistribution of tail mass
@@
    Pattern injection: Adds context-derived markers that create artificial uniformity
@@
    Constant padding: Introduces systematic patterns unrelated to content
@@
    Case flipping: Deterministic transformation that flattens stylistic variation
These transformations preserve semantic content while implementing the key insight from our theoretical analysis: adversaries attack by reducing the "peakiness" of response distributions that our mechanisms rely on for discrimination. The standardization attack particularly mirrors the theoretical construction—both replace natural variation with artificial uniformity.
4.3 Computational Considerations
Our pairwise comparison design generated comprehensive statistical evidence across all conditions. For deployment, all mechanisms require only single API calls per evaluation—MI/DoE and GPPM need only log probabilities without generation, while TVD-MI and LLM judges use one generation call each. We used GPT-4o-mini throughout for consistency and cost management.
@@ 5.1 Information-Theoretic Mechanisms Detect Effectively
All three information theoretic mechanisms successfully discriminate between information preserving and information degrading agents across every tested domain. This validates our theoretical prediction that mechanisms designed for gaming resistance also identify quality differences.
Table 2 presents the discrimination analysis comparing Good Faith agents (Faithful and Style categories) against Problematic agents (Strategic and Low Effort categories). The results starkly contrast two evaluation paradigms:
Information theoretic mechanisms (designed for strategic robustness):
  • 10/10 domains achieve d > 0.5 for all three mechanisms
  • Mean effect sizes: MI (1.87), GPPM (2.70), TVD-MI (5.20)
  • Consistent performance across compression ratios
Direct quality assessment (traditional evaluation):
  • LLM judge without context: 4/9 domains achieve d > 0.5
  • LLM judge with context: 8/9 domains achieve d > 0.5
  • Baseline metrics (ROUGE/BLEU): 5/10 domains achieve d > 0.5
We designed these mechanisms to be incentive compatible (resistant to gaming), yet they outperform methods explicitly designed for quality assessment. This empirical validation in domains with ground truth (where we can verify correlation with ROUGE/BLEU) suggests they will maintain effectiveness in domains without ground truth where robust evaluation is most needed.
Table 2: Effect sizes (Cohen’s d) for discrimination between Good Faith and Problematic agents. Values show mean ± 95% CI. Bold indicates p < 0.001, regular text p < 0.05, gray text non-significant.
@@
SamSum (4.8:1) 0.11 2.52±0.26 2.52±0.23 6.14±0.62 2.70±0.28 0.54±0.15
PubMed (6.7:1) 0.86 2.01±0.33 3.19±0.51 6.53±0.72 8.14±0.93 3.25±0.47
Multi-News (9.0:1) 0.88 1.53±0.18 2.70±0.31 6.55±0.85 4.06±0.59 0.54±0.16
BillSum (9.3:1) 0.91 2.24±0.23 3.59±0.37 5.91±0.74 4.24±0.51 0.16±0.14
CNN/Daily (13.8:1) 0.61 2.06±0.20 3.42±0.36 5.87±0.76 3.55±0.36 0.73±0.10
Reddit TIFU (16.1:1) 0.13 2.52±0.27 3.76±0.36 7.24±0.85 2.70±0.36 0.05±0.14
XSum (18.5:1) 0.29 1.89±0.20 2.85±0.25 6.69±0.68 3.39±0.39 -0.28±0.15

### v2

@@ 3.3 Statistical Limits for Gaming-Resistance
Let $B$ be any distribution-free estimator providing a $(1-\delta)$ confidence lower bound on $I_{f}(X;Y)$ (Def. 3.1), derived from a finite sample empirical type $\mathcal{T}(S^{(N)})$ where $S^{(N)}\sim P_{XY}^{(N)}$. For integers $k\geq 1$ and $N\geq 2$, with probability at least $1-\delta-1/k$ over the sampling:
$B\bigl{(}\mathcal{T}(S^{(N)}),\delta\bigr{)}\leq\frac{1}{2kN^{2}}f(2kN^{2})+\left(1-\frac{1}{2kN^{2}}\right)f(0).$
This bound yields a clean separation between piecewise-linear and super-linear $f$-divergences. For TVD with $f(t)=\tfrac{1}{2}|t-1|$, the ceiling simplifies to $B\leq\ 1-\frac{1}{2kN^{2}}$ so the worst-case certifiable value approaches 1 at rate $1-\Theta(1/N^{2})$. By contrast, for KL with $f(t)=t\log t$, we have $B\leq\ \log(2kN^{2})$ implying that certifying an additional $s$ nats requires $N\geq\Theta(e^{s/2})$ samples in the worst case. Thus bounded, piecewise-linear $f$ admit ceilings that grow polynomially with $N$, whereas unbounded $f$ have ceilings that scale only logarithmically, making per-bit certification exponentially costly under our construction.
@@
Figure 1 (Right) shows the adversarial “mode collapse” construction that drives the bound: keep the largest $kN^{2}$ parts of the response distribution unchanged, spread the next $kN^{2}$ likely responses uniformly at height $1/(2kN^{2})$, and drop the rest. We make this precise by a maximal coupling between the true law $P$ and the surrogate $\widetilde{P}$ that (i) identifies the top $kN^{2}$ atoms, (ii) maps the next $kN^{2}$ atoms to the uniform “orange” cloud, and (iii) annihilates the remainder.
Because $\widetilde{P}$ has only $2kN^{2}$ support points, Lemma F.1 (Maximum MI) implies
$I_{f}(\widetilde{P})\;\leq\;\frac{1}{2kN^{2}}\,f(2kN^{2})+\Bigl{(}1-\frac{1}{2kN^{2}}\Bigr{)}f(0).$
[…]
$\Pr\!\bigl{[}B(\mathcal{T}(S^{(N)}),\delta)>\text{ceiling}\bigr{]}\;\leq\;\delta+\tfrac{1}{k},$
which rearranges to the claimed bound. ∎
This analysis extends McAllester and Stratos (2020) from Shannon information to general $f$-divergences, revealing that robustness depends on the choice of divergence. Showing this generalization required introducing techniques. (i) An explicit coupling that aligns $P$ with a $2kN^{2}$-support surrogate, yielding type-indistinguishability on pure samples; (ii) a maximum MI lemma (Lemma F.1) showing the uniform coupling extremizes $f$-information under support constraints; and (iii) a sharper failure probability of $\delta+\tfrac{1}{k}$ (improving the previous $1.01/k$) via a tight birthday bound within the orange layer.
While Theorem 3.3 considers worst-case mathematical constructions, real adversaries employ semantically plausible attacks. Our experiments (Section 5) test four such strategies. Each approximates the theoretical mode collapse by reducing natural variation while preserving semantic content supporting that our theoretical limits capture practical vulnerabilities.
3.4 Implementing Variational Chain-of-Thought
Computing exact mutual information for high-dimensional text is intractable. Instead, we employ a variational lower bound achievable through categorical classification coupled with a structured chain-of-thought (CoT) reasoning policy for the overseer.
TVD-MI via Type-Based Tests.
Let $P^{+}:=P_{ij}$ denote the joint distribution of paired responses (same source) and $P^{-}:=P_{i}\otimes P_{j}$ denote the product of marginals (independent sources). For total variation distance, the overseer’s reasoning map is a test $r:\mathcal{T}(S)\to\mathcal{C}$ applied to the empirical joint type of a sample $S$. With acceptance set $A$ and f(t)=|t1|f(t)=|t-1|, this yields
$I_{\mathrm{TVD}}(Y_{i};Y_{j})=\mathrm{TV}(P^{+},P^{-})\;\geq\;\hat{I}_{f}^{r}(\mathcal{T}(S)):=\mathrm{TPR}_{r}+\mathrm{TNR}_{r}-1,$ (3)
where now
$\mathrm{TPR}_{r}$ $:=\Pr_{S\sim(P^{+})^{N}}[r(\mathcal{T}(S))\in A],$ (4)
$\mathrm{TNR}_{r}$ $:=\Pr_{S\sim(P^{-})^{N}}[r(\mathcal{T}(S))\not\in A].$ (5)
The bound is tight when $r$ perfectly separates the distributions (Tsybakov, 2008, Definition 2.4). This is an instance of Youden’s (1950) J statistic ($\text{TPR}+\text{TNR}-1$), which measures informativeness (Powers, 2012).
Chain-of-Thought as a Reasoning Strategy.
We understand $r$ as a chain-of-thought program that maps elements of the empirical type to categories through interpretable intermediate steps: (i) salience extraction over the pairs of $\mathcal{T}(S)$, (ii) CoT that provides a contextualized interpretation for the judgment, and (iii) decision via a calibrated thresholding rule.
TVD-MI as a Principled LLM Judge.
Our implementation reveals that TVD-MI can be viewed as an LLM judge with different design choices:
  • Prompt structure: Information relationships ("same source?") vs quality ("which is better?")
  • Aggregation: Information-theoretic (TPR + TNR - 1) vs win-rate averaging
  • Guarantees: DPI-based gaming resistance vs none
Both use identical computational resources (single LLM calls), but our information-theoretic framing provides provable robustness properties.
@@ 4 Study Setup
We designed and pre-registered [^1] Pre-registration available at https://osf.io/c7pum [^2] Code available at https://github.com/zrobertson466920/llm-peer-prediction/tree/main an evaluation study to test whether information-theoretic mechanisms can reliably detect strategic manipulation in AI-generated content. Our study addresses three primary research questions, mapped to our pre-registered hypotheses:
RQ1: Can mechanisms detect agent manipulation strategies? Our method is to use Cohen’s $d$ (standardized mean difference) between Good Faith and Problematic agents. We test H1a ($d>0.5$, medium effect size), H1b (compression effects), H1c (TVD-MI superiority).
RQ2: Do mechanisms produce reliable item-level quality scores? Our method is to calculate item-level AUC (area under ROC curve) for Faithful–Faithful vs. Faithful–Problematic pairs. We test H2c (gaming resistance). We note this was added during analysis as complementary test of pre-registered hypothesis.
RQ3: Do information-theoretic mechanisms resist adversarial attacks? Our method is to measure performance degradation under four tampering strategies. This tests H2a (bounded consistency), H2b (log-prob degradation), H2c (gaming resistance).
Key deviations from pre-registration: (1) Expanded from 3 to 10 domains for proper compression analysis, (2) Collapsed 4 categories to 2 (Good Faith/Problematic) following our theoretical framework, (3) Added AUC analysis recognizing it directly tests gaming resistance. See Appendix A for complete details.
4.1 Experimental Design
Domain Selection. We selected 10 domains spanning compression ratios from 1.1:1 (translation) to 20.2:1 (peer review). See Appendix A.4 for more details. This range tests mechanisms from near-isomorphic tasks to extreme compression where most information is discarded.
Agent Taxonomy. We developed a taxonomy of 29-30 agent strategies per domain, grouped into two categories: good faith agents that preserve information and problematic agents that degrade it. Good faith agents consist of faithful agents (4-6 variants) that are prompted to accurately complete the task and style agents (10-16 variants) that are prompted to preserve information in an alternative presentation. Problematic agents consist of strategic agents (4-10 variants) that are prompted to deliberately manipulate their completions and low effort agents (4-5 variants) that are prompted to give minimal effort or generic responses.
Good Faith agents (Faithful + Style) preserve information while Problematic agents (Strategic + Low Effort) degrade it. Full taxonomy details appear in Appendix C - E.
4.2 Evaluation Mechanisms
We implement three mechanisms with similar computational costs (single API calls):
Information-Theoretic Mechanisms: We evaluate three approaches. The first, MI (DoE), estimates mutual information via the difference of entropies using Llama 3.3-70B log probabilities. The second, GPPM, is the generative peer prediction baseline (Lu et al., 2024). The third, TVD-MI, computes mutual information through total variation distance using a categorical critic (GPT-4o-mini).
Comparison Methods: For baselines, we include LLM Judge (Quality), which uses GPT-4o-mini to assess normative quality via pairwise comparisons. This uses the same model as our TVD-MI critic but prompts for quality judgments rather than information relationships, following Zheng et al. (2023). We also report two standard reference-based metrics: ROUGE, for summarization quality (Lin, 2004), and BLEU, for translation quality (Papineni et al., 2002).
For multiple agents, we extend our framework by aggregating pairwise terms:
$u_{i}:=\sum_{j\neq i}\text{Mechanism-Score}(i,j).$ (6)
This aggregation preserves the gaming-resistance property while enabling scalable implementation across larger agent populations.
@@ 4.2.1 Measurement Tampering Experiments
Our measurement tampering experiments test practical analogues of the adversarial construction from Theorem 3.3, which shows how adversaries degrade estimation by creating more uniform distributions:
@@
    Format standardization: Enforces consistent structure (reduces distributional "peakiness")
@@
    Pattern injection: Adds artificial uniformity via context markers
@@
    Constant padding: Introduces systematic noise patterns
@@
    Case flipping: Flattens stylistic variation deterministically
These preserve semantic content while reducing the natural variation our mechanisms use for discrimination.
4.3 Statistical Analyses
For RQ1 (manipulation detection), we compute paired Cohen’s d between Good Faith and Problematic categories with bootstrap CIs. For RQ2 (decomposability), we analyze item-level AUC distinguishing Faithful-Faithful from Faithful-Problematic pairs. For RQ3 (robustness), we apply four adversarial transformations and measure degradation in both d and AUC.
@@ 5.1 Information-Theoretic Mechanisms Detect Effectively
All three information theoretic mechanisms successfully discriminate between information preserving and information degrading agents across every tested domain. This supports our theoretical prediction that mechanisms designed for gaming resistance also identify quality differences. Table 2 presents the discrimination analysis comparing Good Faith agents (Faithful and Style categories) against Problematic agents (Strategic and Low Effort categories).
For information-theoretic mechanisms designed for strategic robustness, all ten domains achieve $d>0.5$ across the three mechanisms. The mean effect sizes are substantial with MI (1.87), GPPM (2.70), and TVD-MI (5.20). Performance was consistent across different compression ratios. In contrast, direct quality assessment methods show weaker results. Using LLM Judge without context, only six of ten domains surpass $d>0.5$, while with context, nine of ten domains do. Baseline metrics (ROUGE and BLEU) reach this threshold in only six of ten domains.
Table 2: Effect sizes (Cohen’s d) for discrimination between Good Faith and Problematic agents. Cohen’s d measures the standardized difference between group means, with d > 0.5 indicating medium effects and d > 0.8 large effects. Values show mean ± 95% CI. Values show mean ± 95% CI. Bold indicates p < 0.001, regular text p < 0.05, gray text non-significant.
@@
SamSum (4.8:1) 0.11 2.52±0.26 2.52±0.23 6.14±0.64 2.70±0.27 0.54±0.15
PubMed (6.7:1) 0.86 2.01±0.33 3.18±0.50 6.53±0.72 8.14±0.90 3.25±0.47
Multi-News (9.0:1) 0.88 1.53±0.19 2.70±0.30 6.55±0.86 4.06±0.59 0.54±0.16
BillSum (9.3:1) 0.91 2.24±0.24 3.59±0.37 5.91±0.71 4.23±0.48 0.16±0.14
CNN/Daily (13.8:1) 0.61 2.06±0.20 3.42±0.37 5.87±0.76 3.55±0.35 0.72±0.10
Reddit TIFU (16.1:1) 0.13 2.52±0.27 3.76±0.35 7.23±0.84 2.70±0.36 0.05±0.14
XSum (18.5:1) 0.29 1.89±0.19 2.85±0.25 6.69±0.68 3.39±0.39 -0.28±0.14